OpenStudy (anonymous):

(1,2) (3,4) FIND SLOPE

OpenStudy (anonymous):

AND POINT SLOPE FORM

OpenStudy (anonymous):

y-2=1(x-1) <== point-slope

OpenStudy (nenadmatematika):

y-2=1(x-1)

OpenStudy (anonymous):

SOLVE FOR Y

OpenStudy (anonymous):

maybe you can try that one?

OpenStudy (anonymous):

Y=1X +1

OpenStudy (nenadmatematika):

if you want the point slope form you should live the expression I wrote, you wrote slope intercept form